OpenStudy (theraggedydoctor):

For what value of \(x\) is \(\Large\frac{ 2x }{ 3x-1 }\) undefined?

rishavraj (rishavraj):

its undefined when the denominator is 0

OpenStudy (anonymous):

Remove the parentheses around the expression 3x−1. 3x−1=0 Since −1 does not contain the variable to solve for, move it to the right hand side of the equation by adding 1 to both sides. 3x=1 Divide each term in the equation by 3. 3x/3=1/3 Cancel 3 in the numerator and denominator. ( 3 x)/( 3 )=1/3 Divide x by 1 to get x. x=1/3